OpenStudy (anonymous):

How do I use multiplication here to solve this problem? Please help. x(x^2-4)/(x-1)^3 X X-1/X+2

OpenStudy (anonymous):

Is it like this..\[\frac{x(x^2-4)}{(x-1)^3}*\frac{(x-1)}{(x+2)}\]

OpenStudy (anonymous):

@desibaybee

OpenStudy (anonymous):

If so ..\[\frac{x(x-2)(x+2)}{(x-1)^3}*\frac{(x-1)}{(x+2)}\]Now cut off common terms..

OpenStudy (anonymous):

so x(x-2)/(x-1)^2

OpenStudy (anonymous):

@Miyuru

OpenStudy (anonymous):

Yeah..

OpenStudy (anonymous):

Woooooooo! thank u

OpenStudy (anonymous):

@Miyuru

OpenStudy (anonymous):

No problem.. :)

OpenStudy (anonymous):

@Miyuru SO IF IM USING DIVISION FOR THIS PROBLEM: 4X^2/7 / X^2/14

OpenStudy (anonymous):

I SKIP FLIP AND THEN DIVIDE SO 4X^2/7 TIMES 14/X^2 = 4X2/1 SO = 8/1= 8

OpenStudy (anonymous):

So my answer is correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeah that's correct..:)

OpenStudy (anonymous):

WOOOO THANK YOUUUUUUUUUU @Miyuru

OpenStudy (anonymous):

thank You hun @uzumakhi \

OpenStudy (anonymous):

i do nothing here don't thanx me

OpenStudy (anonymous):

lol you told me my answer was correct

OpenStudy (anonymous):

ok then i accept your thanx..................lol

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@Miyuru Don't go yet I need you!

OpenStudy (anonymous):

lol So to add this rational expression 10/3x^2+5/3x^2 I just add the numerators?

OpenStudy (anonymous):

so 15/ 3x^2

OpenStudy (anonymous):

Alright I am here.

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

So that answer is correct too? @Miyuru

OpenStudy (anonymous):

Yeap tht's correct..

OpenStudy (anonymous):

if you have same denominaor you can add them simply

OpenStudy (anonymous):

SO how would I subtract x+1/7x^2 - x-3/7x^2 ?

OpenStudy (anonymous):

@Miyuru can you subtract x+1 - x-3 ?

OpenStudy (anonymous):

Yeah just subtract..

OpenStudy (anonymous):

How do I subtract that can you show me how to do this problem because I know they have the same denominator @Miyuru

OpenStudy (anonymous):

x+1-(x-3) x+1 -x +3 4

OpenStudy (anonymous):

Yeah @uzumakhi is correct..

OpenStudy (anonymous):

thanx @Miyuru

OpenStudy (anonymous):

oh ok I see so why did you change -x-3 too -x+3?

OpenStudy (anonymous):

@uzumakhi

OpenStudy (anonymous):

If I have different denominators like this problem how would I solve this= \[ \frac{ x+2 }{ 2^{2}-9x+10 } +\frac{ 1 }{ x-2 }\]

OpenStudy (anonymous):

thats a 2x^2-9x+10 ^^^^

OpenStudy (anonymous):

Take the lowest common multiple.....

OpenStudy (anonymous):

First factorize \[2x^2-9x+10\]

OpenStudy (anonymous):

@Miyuru how do I do that?

OpenStudy (anonymous):

do it with middle term split

OpenStudy (anonymous):

2x^2 - 5x -4x +10

OpenStudy (anonymous):

As @uzumakhi said

OpenStudy (anonymous):

so (x-4)(x-5)

OpenStudy (anonymous):

\[\frac{ x+2 }{ (x-5)(x-2) }+\frac{ 1 }{ x-2 }\]

OpenStudy (anonymous):

\[\frac{ x+2 }{ (x-4)(x-5) } + \frac{ 1 }{ x-2 }\]

OpenStudy (anonymous):

how did you get x-2?

OpenStudy (anonymous):